# A Novel Density Profile for Isothermal Cores of Dark Matter Halos

### v3

@@ Appendix D Empirical evolutions of parameters in the self-similar solution
$\log{\left(\frac{\rho_{\rm{c,T25}}}{\rho_{\rm{s}}}\right)}=\frac{1}{1-t}\bigg(0.592-25.71\,t+13.02\,t^{2}\\ -4.426\,t^{3}+24.22\,\sqrt{t}-7.363\,\sqrt[3]{t}\,\bigg)^{-1},$ (25)
$\hskip 31.2982pt\frac{r_{\rm{c,T25}}}{r_{\rm{s}}}=-3.110\,t+2.066\,t^{2}-1.149\,t^{3}\\ +2.229\,\sqrt{t}-0.044\,\sqrt[3]{t},$ (26)
$\hskip 31.2982pt\frac{r_{\rm{s,T25}}}{r_{\rm{s}}}=1+12.455\,t-8.785\,t^{2}+0.556\,t^{3}\\ -8.364\,\sqrt{t}+14.706\,\sqrt[3]{t},$ (27)
$\hskip 38.41139ptn_{\rm{T25}}=1-1.978\,t+2.417\,t^{2}-1.293\,t^{3}\\ -3.550\,\sqrt{t}+5.745\,\sqrt[3]{t}.$ (28)
@@ Appendix E Luminosity & heat conductivity
Figure 7: The luminosity profiles of the $10^{8}{\,\rm M_{\odot}}$ halo evolved with $\sigma/m=21.94\,{\rm cm}^{2}\,{\rm g}^{-1}$, calculated at $T\simeq 0.45\,\tau$. The scatter points and dashed lines are derived from the time evolution of the halo, based on interpolation of the simulated profiles and on the corresponding fitted density and reconstructed velocity dispersion using our density model, respectively. The solid lines show the luminosity profiles computed from the velocity dispersion gradient of the four fitted density profiles and the empirical heat conductivity formula (Equation 32) with $\beta=0.75$. Overall, our density profile provides a reliable basis for computing the luminosity profile, both compared to interpolation of the simulation density in the time evolution–based calculation and relative to other estimations of other fitting functions.
@@
Here, $E(r)=3\sigma_{\rm{1D}}^{2}(r)/2+\Phi(r)$, and $r_{M}$ is the radius that encloses the same mass at time $T+\Delta T$ as the enclosed mass within $r$ at time $T$. In practice, we calculate $\mathrm{D}E/\mathrm{D}T$ as the symmetric difference averaged over $T-\Delta T$ and $T+\Delta T$ with $\Delta T\simeq 0.015\,\tau$. Figure 7 shows these profiles, calculated using values of $\sigma_{\rm{1D}}$ and $\Psi$ interpolated from simulation data, as well as those reconstructed from the density profile with our density function. Additionally, we also calculate luminosity as
@@
where typically $\beta=0.75$ for isolated halos [17]. $\rho$ and $\sigma_{\rm{1D}}$ are taken from the density fits and the velocity dispersion reconstructions of the four density profile choices. From Figure 7, we find that the time–evolution–based calculations produce consistent results when using either simulation profile interpolations or our fitted density profile. In the inner regions, however, our profile yields results that are more consistent with the trend observed just outside the core, whereas the simulation-based interpolation suffers from large uncertainties in this regime. The empirical formula for $\kappa$, combined with the reconstructed velocity dispersion gradient, also produces good agreement when using our density profile, particularly in the innermost regions, where a smaller value of $\beta\simeq 0.25$ would provide an exact match to the luminosity profiles.

### v4

@@ Appendix D Empirical evolutions of parameters in the self-similar solution
{multline}
log(ρc,T25ρs) = 11 - t ( 0.592 - 25.71   t + 13.02   t^2
- 4.426   t^3 + 24.22   t - 7.363   3t   )^-1, {multline}      rc,T25rs = - 3.110   t + 2.066   t^2 - 1.149   t^3
+ 2.229   t - 0.044   3t, {multline}      rs,T25rs = 1 + 12.455   t - 8.785   t^2 + 0.556   t^3
- 8.364   t + 14.706   3t, {multline}       n_T25 = 1 - 1.978   t + 2.417   t^2 - 1.293   t^3
- 3.550   t + 5.745   3t.
@@ Appendix E Luminosity & heat conductivity
Figure 7: The luminosity profiles of the $10^{8}{\,\rm M_{\odot}}$ halo evolved with $\sigma/m=21.94\,{\rm cm}^{2}\,{\rm g}^{-1}$, calculated at $T\simeq 0.45\,\tau$. Top: The scatter points and dashed lines are derived from the time evolution of the halo, based on interpolation of the simulated profiles and on the corresponding fitted density and reconstructed velocity dispersion using our density model, respectively. Bottom: The solid lines show the luminosity profiles computed from the velocity dispersion gradient of the four fitted density profiles and the empirical heat conductivity formula (Equation 19) with $\beta=0.75$.
@@
Here, $E(r)=3\sigma_{\rm{1D}}^{2}(r)/2+\Phi(r)$, and $r_{M}$ is the radius that encloses the same mass at time $T+\Delta T$ as the enclosed mass within $r$ at time $T$. In practice, we calculate $\mathrm{D}E/\mathrm{D}T$ as the symmetric difference averaged over $T-\Delta T$ and $T+\Delta T$ with $\Delta T\simeq 0.015\,\tau$. Figure 7 shows these profiles, calculated using values of $\sigma_{\rm{1D}}$ and $\Psi$ interpolated from simulation data, as well as those reconstructed from the fitted density profile.
Additionally, we adopt an alternative approach in which the luminosity is computed directly from the fitted analytical profile
@@
where typically $\beta=0.75$ for isolated halos [17]. $\rho$ and $\sigma_{\rm{1D}}$ are taken from the density fits and the velocity dispersion reconstructions of the four density profile choices. The results of such calculations are also shown in Figure 7. Although the approaches described by Equations 16 and 18 do not necessarily converge, in the inner regions of the halo, the time evolution–based calculation and the reconstructed velocity dispersion gradient approach based on our fitted profile yield fairly similar luminosity profiles.